OpenStudy (pulsified333):

q 0 50 100 150 200 250 300 C'(q) 22 24 23 31 29 31 37 a) If the fixed cost is $11400, use the average of left- and right-hand sums to determine the total cost of producing 150 units. Answer: $ How much would the total cost increase if production were increased one unit, to 151 units?

OpenStudy (pulsified333):

@zepdrix

OpenStudy (pulsified333):

@ganeshie8

ganeshie8 (ganeshie8):

\[C(q)-C(0) ~~=~~ \int\limits_0^q C'(q)\,dq\]

ganeshie8 (ganeshie8):

It is given that fixed cost is $11400 so, \(C(0) = 11400\) : \[C(q)-11400~~=~~ \int\limits_0^q C'(q)\,dq\]

ganeshie8 (ganeshie8):

You can evaluate the right hand side from the given table. Do you know how to approximate the integral when a table is given ?

OpenStudy (pulsified333):

yes i found the average of the right and left hand sum which is 5625

OpenStudy (pulsified333):

@ganeshie8

ganeshie8 (ganeshie8):

I'm getting a different number, can you show ur work how you got 5625

OpenStudy (pulsified333):

Left hand sum= 50(22+24+23)=3450 Right hand sum=50(31+23+24)=3900 And you take the average of the two which is 5625

OpenStudy (pulsified333):

is that right?

ganeshie8 (ganeshie8):

first two lines are correct third line is wrong

ganeshie8 (ganeshie8):

how can average of 3450 and 3900 be 5625 ?

OpenStudy (pulsified333):

damn i just realized that. For some reason my calculations are off

OpenStudy (pulsified333):

it should be 3,675

OpenStudy (pulsified333):

@ganeshie8

ganeshie8 (ganeshie8):

Yes

OpenStudy (pulsified333):

Im just not sure how to incorporate the fix cost

ganeshie8 (ganeshie8):

fixed is the cost when the number of items is 0 : C(0) is the fixed cost

ganeshie8 (ganeshie8):

\[C(q)-11400~~=~~ \int\limits_0^q C'(q)\,dq\] \[C(150)-11400~~=~~ \int\limits_0^{150} C'(q)\,dq\approx 3675\]

ganeshie8 (ganeshie8):

\[C(150) - 11400 = 3675\] solve \(C(150)\)

OpenStudy (pulsified333):

Oh okay

OpenStudy (pulsified333):

C(150)=15,075

OpenStudy (pulsified333):

@ganeshie8 how do you find the second part?

ganeshie8 (ganeshie8):

use the tangent approximation

ganeshie8 (ganeshie8):

f(x+1) = f(x) + f'(x)*1 = f(x) + f'(x)

ganeshie8 (ganeshie8):

C(150+1) = C(150) + C'(150)

ganeshie8 (ganeshie8):

you have just worked the value of C(150) and you can lookup C'(150) from the given table

OpenStudy (pulsified333):

31+15075?

ganeshie8 (ganeshie8):

Yep!

OpenStudy (pulsified333):

it says its wrong

OpenStudy (pulsified333):

oh never mind i got it because it was asking for the amount it increased which was 31:D